# Robust methods for Darcy–Forchheimer equationsDas, Hutridurga, Pani, Ruiz-Baier

### v1

@@ 4.2 A priori error analysis
The proof follows directly from the quasi-optimality estimate (4.10) and the above approximation properties, in the following manner
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}}+\|p-p_{h}\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ $\leq{C_{c}}\bigl(\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})+\mathrm{dist}(p,\mathrm{Q}_{h})\bigr)$
${\leq{\widetilde{C}_{c}}\bigl(\|\boldsymbol{u}-{\widetilde{\Pi}_{h}}(\boldsymbol{u})\|_{\mathbf{V}}+\|p-{\widetilde{P}_{h}}(p)\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}\bigr)}$
$\leq C_{\flat}h^{l+1}\bigl\{\kappa^{-\frac{1}{2}}\|\boldsymbol{u}\|_{l+1,\Omega}+\|{\mathop{\mathrm{div}}\nolimits}\,\boldsymbol{u}\|_{l+1,\Omega}+\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{u}\|_{l+1,3,\Omega}\bigr\}$
$\ +C_{\flat}h^{l+1}\{|p|^{2}_{l+1,\Omega}+\kappa|p|^{2}_{l+2,\Omega}+\mathtt{F}^{-\frac{2}{3}}|p|_{l+2,\frac{3}{2},\Omega}\}^{\frac{1}{2}},$
where, $\widetilde{\Pi}_{h},\widetilde{P}_{h}$ are scaled (with suitable constants independent of $h,\kappa,\mathtt{F}$) counterparts of $\Pi_{h}$ and $P_{h}$, and $C_{\flat}=\widetilde{C}_{c}\max\{\acute{C},C_{\sharp}\}$. This completes the rest of the proof.
@@ 5 Preconditioning for the linearised Darcy–Forchheimer equations
$\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$ $:=\bigl[\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap\{\mathtt{F}\mathcal{H}(\hat{\boldsymbol{u}})\}^{\frac{1}{3}}\mathbf{L}^{3}(\Omega)\bigr]$
(5.2) $\quad\times\bigl[\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}_{\Gamma_{p}}^{1}(\Omega)+\{\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}\}^{-\frac{1}{3}}\mathrm{W}_{\Gamma_{p}}^{1,\frac{3}{2}}(\Omega)\bigr],$
@@
Assume that for any $\boldsymbol{w}\in\underline{\mathbf{V}},$ the function $\mathtt{F}\mathcal{H}(\boldsymbol{w})$ is in $\mathbf{L}^{\infty}(\Omega)$ and that $\mathtt{F}\mathcal{H}(\boldsymbol{w})(\boldsymbol{x})\geq 0$ a.e. in $\Omega$. Then, the problem (5.3) has a unique solution in $\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$.
@@
is bounded in $\underline{\mathbf{V}}$ using the Hölder’s inequality and coercive in $\underline{\mathbf{Z}}=\mathrm{ker}B$; all with constants independent of $\kappa$, $\mathtt{F}$, and $\mathcal{H}(\boldsymbol{u}^{m}).$ Finally, the boundedness and inf-sup condition on $B$ have been shown in the proof of Theorem 3.4 for a different weight. In this case, we only need to alter the scaling, which completes the rest of the proof.
@@
Let $(\hat{\boldsymbol{u}},\hat{p})\in\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$ be the state around which the Newton–Raphson linearisation is performed. Then the following block diagonal operator
(5.4) $\begin{pmatrix}\mathcal{P}:=(\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}{\mathcal{I}}-\nabla\mathop{\mathrm{div}}\nolimits)^{-1}&0\\ 0&\!\!\!\!\!\!\!\!\!\!\!\mathcal{Q}:=\mathcal{I}^{-1}+{\bigl(-\kappa\Delta\bigr)^{-1}-\bigl([\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta_{\frac{3}{2}}}\bigr)^{-1}}\end{pmatrix}$,
is a parameter-robust preconditioner for (5.3), in the sense that the condition numbers of the preconditioned matrix are uniformly bounded in $\kappa$ and $\mathtt{F}$. Here, making abuse of notation, by $\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta_{\frac{3}{2}}}$ we denote the (singular) weighted p-Laplace operator associated with the space $\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\Omega)$, defined as
$\langle[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}\Delta_{\frac{3}{2}}\phi,\psi\rangle:=\int_{\Omega}[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}|\nabla\phi|^{-\frac{1}{2}}\nabla\phi\cdot\nabla\psi,\qquad\forall\psi\in\mathrm{W}^{1,3}(\Omega).$
@@
(5.5) $\|q\|^{2}_{\underline{\mathrm{Q}}}:=\inf_{\varphi\in\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)}\bigl\{\|q-\varphi\|^{2}_{0,\Omega}+\kappa\|\nabla\varphi\|^{2}_{0,\Omega}+{[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}}\|{\nabla}\varphi\|^{\frac{3}{2}}_{0,\frac{3}{2},\Omega}\bigr\}.$
Note that the above infimum is attained by a $\varphi\in\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)$ (therefore also in $\mathrm{L}^{2}(\Omega)$ and $\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\Omega)$) that satisfies
(5.6a) $(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{{\Delta}_{\frac{3}{2}}})\varphi$ $=q\quad\text{in }\Omega,$
(5.6b) $(\kappa\nabla\varphi+[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}|\nabla\varphi|^{-\frac{1}{2}}\nabla\varphi)\cdot\boldsymbol{n}$ $=0\quad\text{on }\Gamma_{\boldsymbol{u}},$
@@
The unique solvability of (5.6) follows from the Minty–Browder framework for monotone operators, stressing that $q$ is from a suitable dual space (following from Sobolev embedding). Further details are omitted. Using (5.6), we obtain the following
$\|q\|^{2}_{\underline{\mathrm{Q}}}$ $=\|q-\varphi\|^{2}_{0,\Omega}+\kappa\|\nabla\varphi\|^{2}_{0,\Omega}+[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}\|{\nabla}\varphi\|^{\frac{3}{2}}_{0,\frac{3}{2},\Omega}$
$=\langle-\kappa\Delta\varphi-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}}\varphi,-\kappa\Delta\varphi-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}}\varphi\rangle-\kappa\langle\Delta\varphi,\varphi\rangle$
$\qquad-\langle[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}}\varphi,\varphi\rangle$
$=\langle(-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}})\varphi,(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}})\varphi\rangle.$
@@
(5.7) $\|q\|^{2}_{\underline{\mathrm{Q}}}=\langle(-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}})(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}})^{-1}q,q\rangle.$
@@
$\mathcal{Q}=(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}{\Delta}_{\frac{3}{2}})^{-1}.$
Therefore, proceeding similarly as in [28, Section 3], we can assert that (5.4) is a $(\kappa,\mathtt{F})$-robust block diagonal preconditioner for system (5.3). This concludes the proof.
@@
$\underline{\underline{\mathbf{V}}}\times\underline{\underline{\mathrm{Q}}}:=(\kappa^{-1}+\mathtt{F}\mathcal{H}(\hat{\boldsymbol{u}}))[\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)]\times(\kappa^{-1}+\mathtt{F}\mathcal{H}(\hat{\boldsymbol{u}}))^{-1}\,\mathrm{L}^{2}(\Omega),$
where again $(\hat{\boldsymbol{u}},\hat{p})\in\underline{\underline{\mathbf{V}}}\times\underline{\underline{\mathrm{Q}}}$ is the state around which the Newton–Raphson linearisation is performed. This space is associated with the block diagonal Riesz map
(5.8) $\begin{pmatrix}\mathcal{P}:=\bigl([\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]\,({\mathcal{I}}-\nabla\mathop{\mathrm{div}}\nolimits)\bigr)^{-1}&0\\ 0&\!\!\!\!\!\!\!\!\!\!\!\mathcal{Q}:=([\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}(\hat{\boldsymbol{u}})}]^{-1}\,\mathcal{I})^{-1}\end{pmatrix}$.
The parameter-robustness for this preconditioner can be easily inferred from a scaling argument as in [28, Example 3.2]. Nevertheless this preconditioner might not be of much interest in the case of coupling with other multiphysics effects. This has been discussed in length in [5].

### v2

@@ 4.2 A priori error analysis
We first take $m=0$ in (4.19a), leading to
$\|\boldsymbol{u}-\Pi_{h}\boldsymbol{u}\|_{0,t,\Omega}\leq C_{1}h^{l}|\boldsymbol{u}|_{l,t,\Omega},$
and we use this bound with $t\in\{2,r\}$. We also take $m=0$ and $t=2$ in (4.19b), which gives
$\|\mathop{\mathrm{div}}\nolimits(\boldsymbol{u}-\Pi_{h}\boldsymbol{u})\|_{0,\Omega}\leq C_{2}h^{l}|\boldsymbol{u}|_{l,\Omega},$
and select $1\leq l\leq k+1$ for the last two displayed estimates. Therefore, owing to Corollary 4.10, we arrive at the velocity error bound
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}}\leq C_{D}\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})\leq C_{D}C_{1}^{2}C_{2}h^{l}(\kappa^{-\frac{1}{2}}\|\boldsymbol{u}\|_{l,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{u}\|_{l,\Omega}+\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{u}\|_{l,r,\Omega}).$
For the pressure we employ the decomposition $p=p_{1}+p_{2}+p_{3}$, the properties of the sum norm (4.4), apply (4.19c) with $m=0$ and $t=2$, and (4.19d) once with $t=2$, and once with $t=r^{\prime}$, both summed over $K\in\mathcal{T}_{h}$. We combine that with the following trace inequality to control the jump term in the broken norms
$h_{F}^{1-t}\|[\![p_{j}-P_{h}p_{j}]\!]\|^{t}_{0,t,F}=h_{F}^{1-t}\|[\![P_{h}p_{j}]\!]\|^{t}_{0,t,F}\leq C_{4}h^{lt}|p_{j}|^{t}_{l+1,r,\omega_{F}},$
where $\omega_{F}$ stands for the usual face patch. We apply this for $j=2,3$ with $t=2$ and $t=r^{\prime}$), and we readily obtain
$\|p-P_{h}p\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ $\leq\|p_{1}-P_{h}p_{1}\|_{0,\Omega}+\kappa^{\frac{1}{2}}|p_{2}-P_{h}p_{2}|_{1,\mathcal{T}_{h}}+\mathtt{F}^{-\frac{1}{r}}|p_{3}-P_{h}p_{3}|_{1,r^{\prime},\mathcal{T}_{h}}$
$\leq\max\{C_{2},C_{3},C_{4}\}h^{l}\bigl(|p_{1}|_{l,\Omega}+\kappa^{\frac{1}{2}}|p_{2}|_{l+1,\Omega}+\mathtt{F}^{-\frac{1}{r}}|p_{3}|_{l+1,r^{\prime},\Omega}\bigr).$
Then we can invoke directly the quasi-optimality estimate (4.11), to get the desired result with $C_{\sharp}=C_{D}(2C_{1}+C_{2})$ and $C_{\flat}=\widetilde{C}_{g}\max\{C_{2},C_{3},C_{4}\}$. Since the last two contributions of the pressure norm require derivatives, the $\mathrm{L}^{2}$ projection only delivers theoretical approximability if $k\geq 1$ or if the infimum decomposition can be chosen such that $p=p_{1}$ (and $p_{2}=p_{3}=0$), i.e., if the optimal decomposition lives entirely in the $\mathrm{L}^{2}$ component.
@@ 5 Preconditioning for the linearised Darcy–Forchheimer equations
$\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$ $:=\bigl[\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap{[\mathtt{F}\mathcal{H}({\boldsymbol{u}^{m}})]^{\frac{1}{r}}\mathbf{L}^{r}(\Omega)}\bigr]$
(5.2) $\quad\times\bigl[\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}_{\Gamma_{p}}^{1}(\Omega)+{[\mathtt{F}\mathcal{H}({\boldsymbol{u}^{m}})]^{-\frac{1}{r}}\mathrm{W}_{\Gamma_{p}}^{1,r^{\prime}}(\Omega)}\bigr],$
@@
Assume that for any $\boldsymbol{w}\in\underline{\mathbf{V}},$ the function $\mathtt{F}\mathcal{H}(\boldsymbol{w})$ is in $\mathbf{L}^{\infty}(\Omega)$ and that $\mathtt{F}\mathcal{H}(\boldsymbol{w})(\boldsymbol{x})\geq C>0$ a.e. in $\Omega$. Then, the problem (5.3) has a unique solution in $\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$.
@@
is bounded in $\underline{\mathbf{V}}$ using the Hölder’s inequality and coercive in $\underline{\mathbf{Z}}=\mathrm{ker}B$; all with constants independent of $\kappa$, $\mathtt{F}$, and $\mathcal{H}(\boldsymbol{u}^{m}).$ Finally, the boundedness and inf-sup condition on $B$ have been shown in the proof of Theorem 3.4 for a different weight. In this case, we only need to alter the scaling, thus completing the proof.
Under the assumption in Lemma 5.1, the weighted $r$-Laplacian defines a coercive and bounded operator on $\mathrm{W}^{1,r}_{\Gamma_{p}}(\Omega)$ with constants independent of the discretisation parameters. From a practical standpoint, this corresponds to freezing the nonlinear coefficient at the current iterate and using a uniformly elliptic linear operator at each step. In all numerical experiments reported in Section 6, the iterates remain in a regime where these bounds are satisfied, and no degeneracy or loss of stability is observed.
@@
Let $({\boldsymbol{u}^{m}},{p^{m}})\in\underline{\mathbf{V}}\times\underline{\mathrm{Q}}$ be the state around which the Newton–Raphson linearisation is performed. Then the following block diagonal operator
(5.4) $\begin{pmatrix}\mathcal{P}:=(\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}{\mathcal{I}}-\nabla\mathop{\mathrm{div}}\nolimits)^{-1}&0\\ 0&\!\!\!\!\!\!\!\!\!\!\!\mathcal{Q}:=\mathcal{I}^{-1}+{\bigl(-\kappa\Delta\bigr)^{-1}-\bigl([\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta_{{r^{\prime}}}}\bigr)^{-1}}\end{pmatrix}$,
is a parameter-robust preconditioner for (5.3), in the sense that the condition numbers of the preconditioned matrix are uniformly bounded in $\kappa$ and $\mathtt{F}$. Here, making abuse of notation, by $\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta_{r^{\prime}}}$ we denote the (singular) weighted p-Laplace operator associated with the space $\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\Omega)$, defined as
$\langle[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}\Delta_{r^{\prime}}\phi,\psi\rangle:=\int_{\Omega}[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}|\nabla\phi|^{r^{\prime}-2}\nabla\phi\cdot\nabla\psi,\qquad\forall\psi\in\mathrm{W}^{1,r}(\Omega).$}
@@
(5.5) $\|q\|^{2}_{\underline{\mathrm{Q}}}:=\inf_{\varphi\in\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)}\bigl\{\|q-\varphi\|^{2}_{0,\Omega}+\kappa\|\nabla\varphi\|^{2}_{0,\Omega}+{{[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}}\|{\nabla}\varphi\|^{r^{\prime}}_{0,r^{\prime},\Omega}}\bigr\}.$
Note that the above infimum is attained by a $\varphi\in\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)$ (therefore also in $\mathrm{L}^{2}(\Omega)$ and $\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\Omega)$) that satisfies
(5.6a) $(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{{\Delta}_{r^{\prime}}})\varphi$ $=q\quad\text{in }\Omega,$
(5.6b) $(\kappa\nabla\varphi+[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}|\nabla\varphi|^{r^{\prime}-2}\nabla\varphi)\cdot\boldsymbol{n}$ $=0\quad\text{on }\Gamma_{\boldsymbol{u}},$
@@
The unique solvability of (5.6) follows from the Minty–Browder framework for monotone operators, stressing that $q$ is from a suitable dual space (following from Sobolev embedding). Using (5.6), we obtain the following
$\|q\|^{2}_{\underline{\mathrm{Q}}}$ $=\|q-\varphi\|^{2}_{0,\Omega}+\kappa\|\nabla\varphi\|^{2}_{0,\Omega}+[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}\|{\nabla}\varphi\|^{r^{\prime}}_{0,r^{\prime},\Omega}$
$=\langle-\kappa\Delta\varphi-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}}\varphi,-\kappa\Delta\varphi-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}}\varphi\rangle-\kappa\langle\Delta\varphi,\varphi\rangle$
$\qquad-\langle[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}}\varphi,\varphi\rangle$
$=\langle(-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}})\varphi,(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}})\varphi\rangle.$
@@
(5.7) $\|q\|^{2}_{\underline{\mathrm{Q}}}={\langle(-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}})(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}})^{-1}q,q\rangle}.$
@@
$\mathcal{Q}=(\mathcal{I}-\kappa\Delta-[\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{r-4}{\Delta}_{r^{\prime}})^{-1}$}.
Therefore, proceeding similarly as in [29, Section 3], we can assert that (5.4) is a $(\kappa,\mathtt{F})$-robust block diagonal preconditioner for system (5.3). This concludes the proof.
The realisation of $\mathcal{Q}$ in (5.4) requires solving a nonlinear $r^{\prime}$-Laplacian problem, which would indeed be computationally expensive. In practice, we do not apply the nonlinear inverse. Instead, the nonlinear coefficient is frozen at the current iterate, yielding a linear, weighted elliptic operator. The application of $\mathcal{Q}^{-1}$ is therefore reduced to the solution of a linear problem with variable coefficients. In the present implementation, this linear system is solved using a sparse direct solver. However, the formulation is fully compatible with optimal-complexity solvers.
@@
$\underline{\underline{\mathbf{V}}}\times\underline{\underline{\mathrm{Q}}}:=(\kappa^{-1}+\mathtt{F}\mathcal{H}({\boldsymbol{u}^{m}}))[\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)]\times(\kappa^{-1}+\mathtt{F}\mathcal{H}({\boldsymbol{u}^{m}}))^{-1}\,\mathrm{L}^{2}(\Omega),$
where again $({\boldsymbol{u}^{m}},{p^{m}})\in\underline{\underline{\mathbf{V}}}\times\underline{\underline{\mathrm{Q}}}$ is the state around which the Newton–Raphson linearisation is performed. This space is associated with the block diagonal Riesz map
(5.8) $\begin{pmatrix}\mathcal{P}:=\bigl([\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]\,({\mathcal{I}}-\nabla\mathop{\mathrm{div}}\nolimits)\bigr)^{-1}&0\\ 0&\!\!\!\!\!\!\!\!\!\!\!\mathcal{Q}:=([\kappa^{-1}\mathcal{I}+\mathtt{F}{\mathcal{H}({\boldsymbol{u}^{m}})}]^{-1}\,\mathcal{I})^{-1}\end{pmatrix}$.
The parameter-robustness for this preconditioner can be easily inferred from a scaling argument as in [29, Example 3.2]. Nevertheless this preconditioner might not be of much interest in the case of coupling with other multiphysics effects. This has been discussed in length in [5]. Note that while in the Darcy limit the operator preconditioning framework we use coincides with that of [5], unfortunately we do not see a clear way to continuously extend our construction to include the case of zero Forchheimer coefficient. This is because the scaling with negative powers of this parameter are hard-coded in the norm and preconditioning terms, and the linear algebra breaks down for $\mathtt{F}=0$. However, we can take positive, arbitrarily small values of this parameter (see, for instance, the behaviour for $\mathtt{F}=10^{-9}$ in Example 4 of Section 6).